OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = 3x + 2, what is f(g(x)) ? 12x + 9 12x + 7 12x + 5 12x + 3

OpenStudy (anonymous):

wherever you see x, replace with g(x)

OpenStudy (anonymous):

\[fog(x)=4(3x+2)+1=12x+8+1=12x+9\]

OpenStudy (anonymous):

in otherwords you're doing f(3x+2)

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

12x+9